Your toaster has a power cord with a resistance of
Answer #1

Resultant Resistance of circuit = 0.022 + 9.9 = 9.922 ohm

potential difference = 120 V

current in circuit = V/R = 120/9.922 = 12.09 A

power dissipated in cord = I^2*R = 12.09^2*0.022 = 3.21 watt

B)

power dissipated in heating element = 12.09^2*9.9 = 1447.06 watt = 1.45 kwatt